Find the remainder when 7^21+49^21+343^21+2401^21 is divided by 7^20+1


1
Expert's answer
2021-06-03T08:12:38-0400

1) "7^{21} + 49^{21} + 343^{21} + 2401{21} ="

"7^{21} + (7\cdot7)^{21}+(7\cdot7\cdot7)^{21} + (7\cdot7\cdot7\cdot7)^{21} ="

"7^{21} +{(7^{21})}^2+{(7^{21})}^3 + {(7^{21})}^4;"

2) "7^{21} = 7\cdot(7^{20})=7\cdot({7^{20} + 1})-7"

3) "Rem[7^{21} \/ (7^{20} + 1)] = -7"

We have represented 721 as 7⋅(720+1)−7 (look step 2)

7⋅(720+1)−7 mod (720+1) = -7 because


"\frac{7\cdot(7^{21}+1) -7}{7^{21}+1} = \frac{7\cdot(7^{21}+1)}{7^{21}+1}+\frac{-7}{7^{21}+1} = 7+\frac{-7}{7^{21}+1}"


where is -7 is our reminder.

For steps 4, 5, 6 We have to raise our reminder to a power.

4) "Rem[{(7^{21})}^2\/(7^{20} + 1)] = (-7)^{2}"

5) "Rem[{(7^{21})}^3\/(7^{20} + 1)] = (-7)^3"

6) "Rem[{(7^{21})}^4\/(7^{20} + 1)] = (-7)^4"

"Answer: -7+(-7)^2+(-7)^3+(-7)^4 = 2100"